Spiegelungssatz に関する文献 (2012 年 3 月更新版)

[1] A. Scholz : Über die Beziehung der Klassenzahlen quadratischer Körper zueinander, J. Reine Angew. Math. 166 (1932), 201–203; doi: 10.1515/crll.1932.166.201 [PDF (624 KB)];

PPN243919689_0166_log20.pdf (435,63 KB), GDZ; [PDF],

DigiZeitschriften; Zbl 0004.05104 (reviewed by Deuring) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 58.0181.05 (reviewed by Reichardt, H.) [html, DVI, ps, PDF, TeX].

[2] H. Reichardt : Arithmetische Theorie der kubischen Körper als Radikalkörper, Monatsh. Math. Phys. 40 (1933), no. 1, 323–350; DOI: 10.1007/BF01708874 [pj383437x2437217.pdf (1.1 MB)]; MR1550211; Zbl 0008.10302 (reviewed by Taussky) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; JFM 59.0195.01 (reviewed by Schwarz, L.) [html, DVI, ps, PDF, TeX].

[3] H.-W. Leopoldt : Zur Struktur der l-Klassengruppe galoisscher Zahlkörper, J. Reine Angew. Math. 199 (1958), 165–174; doi: 10.1515/crll.1958.199.165 [PDF (1274 KB)];

PPN243919689_0199_log13.pdf (1.014,45 KB), GDZ; [PDF],

DigiZeitschriften; MR0096633 (20 #3116) (reviewed by Y. Kawada; reprinted in Reviews in Number Theory, volume 5, R26-12) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0082.25402 (reviewed by Z. Suetuna) [scanned image].

[4] K. Shiratani : Bemerkung zur Theorie der Kreiskörper, Mem. Fac. Sci. Kyushu Univ. Ser. A Math. [九州大學理學部紀要 Ser. A (数学)] 18 (1964), no. 2, 121–126; doi:10.2206/kyushumfs.18.121, JOI:JST.JSTAGE/kyushumfs/18.121 [PDF (388 K)]; MR0172869 (30 #3085) (reviewed by O. F. G. Shilling; reprinted in Reviews in Number Theory, volume 5, R18-45) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0147.03504 (reviewed by E. Maus) [scanned image].

[5] S.-N. Kuroda : Über den allgemeinen Spiegelungssatz für Galoissche Zahlkörper, J. Number Theory 2 (1970), no. 3, 282–297; doi:10.1016/0022-314X(70)90056-9 [PDF (849 K)]; MR0311624 (47 #186) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5A, R26-208) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0222.12013 (reviewed by H. Yokoi) [scanned image].

[6] G. Gras : Extensions abéliennes non ramifiées de degré premier d’un corps quadratique, Bull. Math. Soc.

France 100 (1972), 177–193; BSMF_1972__100__177_0.djvu,
BSMF_1972__100__177_0.pdf, Numdam; MR0302604 (46

#1748) (reviewed by H. W. Brinkmann; reprinted in Reviews in Number Theory, volume 5A, R20-203) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0249.12002 (reviewed by O. Neumann) [scanned image].

[7] S. Yamamoto : On the rank of the p-divisor class group of Galois extensions of algebraic number fields, Kumamoto J. Sci. (Math.) 9 (1972), no. 1, 33–40; MR0302613 (46 #1757) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5A, R26-205) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0252.12009 (Zusammenfassung des Autors) [scanned image].

[8] A. Kudo : On the reflection theorem in prime cyclotomic fields, Mem. Fac. Sci. Kyushu Univ. Ser. A Math. [九州大學理學部紀要 Ser. A (数学)] 26 (1972), no. 2, 333–337; doi:10.2206/kyushumfs.26.333, JOI:JST.JSTAGE/kyushumfs/26.333 [PDF (333K)]; MR0409407 (53 #13162) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5A, R26-231) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0252.12007 (reviewed by H. Koch) [scanned image].

[9] B. Oriat : Relation entre les 2-groupes des classes d’idéaux au sens ordinaire et restreint de certains corps de nombres, Bull. Soc. Math. France 104

(1976), fasc. 3, 301–307; BSMF_1976__104__301_0.djvu,
BSMF_1976__104__301_0.pdf, Numdam, MR0435038 (55

#8000) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5A, R24-223) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0352.12007 (reviewed by H. Yokoi) [scanned image].

[10] B. Oriat : Spiegelungssatz, Séminaire de Théorie des Nombres, Années 1975–1976, 20 pp., Publ. Math. Fac. Sci. Besançon, Univ. Franche-Comté, Besançon, 1976; Oriat1.pdf.

[11] B. Oriat : Relation entre les 2-groupes des classes d’idéaux des extensions quadratiques k(\sqrt{d}) et k(\sqrt{−d}), Ann. Inst. Fourier (Grenoble) 27 (1977), Fasc.

2, vii, 37–59; AIF_1977__27_2_37_0.pdf, Cedram;
AIF_1977__27_2_37_0.djvu,  AIF_1977__27_2_37_0.pdf,

Numdam; MR0466064 (57 #5947) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5A, R14-299) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0351.12001 (Autorreferat) [scanned image].

[12] B. Oriat : Relation entre les 2-groupes des classes d’idéaux de k(\sqrt{d}) et k(\sqrt{−d}), Journées Arithmétiques de Caen (Univ. Caen, Caen, 1976), Astérisque No. 41–42, Soc. Math. France, Paris, 1977, pp. 247–249; MR0506162 (58 #22013)(reprinted in Reviews in Number Theory, volume 5A, R14-299) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0348.12007 (Résumé) [scanned image].

[13] Ph. Satgé : Inégalités de miroir, Séminaire Delange-Pisot-Poitou, 18e Année: 1976/77, Théorie des Nombres, Fasc. 2, Exp. No. 18, 4 pp., Secrétariat

Math., Paris, 1977; SDPP_1976-1977__18_2_A2_0.djvu,
SDPP_1976-1977__18_2_A2_0.pdf, Numdam; MR0551345

(80k:12014) (author’s introduction and note added in proof; reprinted in Reviews in Number Theory, volume 5A, R12-217) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0373.12005 (reviewed by E. Brown) [scanned image].

[14] B. Oriat : Sur la divisibilite par 8 et 16 des nombres de classes d’ideaux des corps quadratiques Q(\sqrt{2p}) et Q(\sqrt{−2p}), J. Math. Soc. Japan [日本数学会欧文誌] 30 (1978), No. 2, 279–285; doi:10.2969/jmsj/03020279, euclid.jmsj/1240432556 [PDF (603 KB)]; JOI:JST. Journalarchive/jmath1948/30.279 [30_279.pdf (494 K)], Journal@rchive, JST; MR0498482 (58 #16592) (reviewed by D. Coray; reprinted in Reviews in Number Theory, volume 5A, R14-299E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0391.12003 (reviewed by H. J. Godwin) [scanned image].

[15] B. Oriat et Ph. Satgé : Un essai de généralisation du “Spiegelungssatz”, J. Reine Angew. Math. 307/308 (1979), 134–159; doi: 10.1515/crll.1979.307-308.134 [PDF (3438

KB)]; PPN243919689_0307_0308_log10.pdf (3,04 MB),

GDZ; [PDF], DigiZeitschriften; MR0534216 (80f:12007) (reviewed by J. Browkin; reprinted in Reviews in Number Theory, volume 5A, R12-217E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0395.12015 (Autorreferat) [scanned image].

[16] B. Oriat : Généralisation du “Spiegelungssatz”, Astérisque 61 (1979), 169–175; Zbl 0403.12014 (reviewed by G. Gras) [scanned image].

[17] Ph. Dutarte : Compatibilité avec le Spiegelungssatz de probabilités conjecturales sur le p-rang du groupe des classes, Théorie Nombres (Besançon), 1983/1984, Exp. No. 4, Publ. Math. Fac. Sci. Besançon, Univ. Franche-Compté, Besançon, 1984, 11 pp.; Dutarte.pdf; MR0803700 (86m:11103) (reviewed by D. A. Buell; reprinted in Reviews in Number Theory, volume 5B, Y40-614) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0576.12004 (reviewed by G. Gras) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[18] F. Diaz y Diaz, P. Llorente and J. Quer : Cubic fields, a congruential criterion for Scholz’s theorem and new real quadratic fields with 3-rank equal to 4, Arch. Math. (Basel) 50 (1988), No. 4, 356–359; DOI: 10.1007/BF01190231 [x80211633v67053p.pdf (216.7 KB)]; MR0937339 (89d:11097) (reviewed by F. Halter-Koch; reprinted in Reviews in Number Theory, volume 5B, R29-741) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0642.12008 (reviewed by K. Lakkis) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[19] J.-F. Jaulent : Dualité dans les corps sur circulaires, Séminaire de Théorie des Nombres, Paris 1986–87 (C. Goldstein, ed.), Progr. Math., vol. 75, Birkhäuser Boston, Boston MA, 1988, pp. 183–220; MR0990512 (90g:11146) (reviewed by L. Washington; reprinted in Reviews in Number Theory, volume 5B, R18-631) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0679.12007 (reviewed by J.-F. Jaulent) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[20] J. Brinkhuis : Normal integral bases and the Spiegelungssatz of Scholz, Acta Arith. LXIX (1995), fasc. 1, 1–9; aa6911.pdf, icm; MR1310838 (96a:11124) (reviewed by J.-F. Jaulent; reprinted in Reviews in Number Theory, volume 5B, R33-753) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0838.11073 (reviewed by J.-F. Jaulent) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[21] D. A. Buell and V. Ennola : On a parameterized family of quadratic and cubic fields, J. Number Theory 54 (1995), no. 1, 134–148; doi:10.1006/jnth.1995.1106 [PDF (491 K)]; MR1352641 (96h:11109) (reviewed by M. Olivier: reprinted in Reviews in Number Theory, volume 5B, R16-715) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0832.11039 (reviewed by R. Mollin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[22] G. Gras : Théorèmes de réflexion, J. Théor. Nombres Bordeaux 10 (1998), Fasc. 2, 399–499; doi: 10.5802/jtnb.2

[JTNB_1998__10_2_399_0.pdf];  JTNB_1998__10_2_
399_0.djvu, JTNB_1998__10_2_399_0.pdf, Numdam;

MR1828251 (2002g:11154) (reviewed by J.-F. Jaulent) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0949.11058 (reviewed by J.-F. Jaulent) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[23] 岸康弘 : Spiegelung-relations between 3-ranks of absolute ideal class groups and congruent ones modulo 3^{2} in quadratic fields, 研究集会報告集 X, 整数論 (早稲田大学, 1999) (足立恒雄編), 早稲田大学, 1999, pp. 7–14.

[24] Y. Kishi : A constructive approach to Spiegelung relations between 3-ranks of absolute ideal class groups and congruent ones modulo (3)^{2} in quadratic fields, J. Number Theory 83 (2000), no. 1, 1–49; doi:10.1006/jnth.1999.2498 [PDF (292 K)]; MR1767650 (2001h:11137) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0967.11048 (reviewed by F. Soriano-Gafiuk) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[25] G. Gras : Ramifications minimales, J. Théor. Nombres Bordeaux 12 (2000), Fasc. 2, 423–435; doi: 10.5802/jtnb.288

[JTNB_2000__12_2_423_0.pdf]; JTNB_2000__12_2_423_0.
djvu, JTNB_2000__12_2_423_0.pdf, Numdam; Gras.ps,

emis; MR1823194 (2002c:11151) (reviewed by J.-F. Jaulent) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0985.11058 (reviewed by F. Soriano-Gafiuk) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[26] R. Sujatha : Index of the Stickelberger ideal and a reflection theorem, Cyclotomic fields and related topics (Pune, 1999) (S. D. Adhikari, S. A. Katre, D. Thakur, eds.), Bhaskaracharya Pratishthana, Pune, 2000, pp. 155–164; MR1802382 (2002a:11120) (reviewed by R. Kučera) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1038.11070 (reviewed by R. Kučera) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[27] Y. Lee : Cohen-Lenstra heuristics and Spiegelungssatz, J. Number Theory 92 (2002), no. 1, 37–66; doi:10.1006/jnth.2001.2699 [PDF (222 K)]; MR1880583 (2002j:11130) (reviewed by G.-D. Villa-Salvador) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1026.11080 (reviewed by K. Yamamura) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[28] Y. Kishi : The Spiegelungssatz for p = 5 from a constructive approach, Math. J. Okayama Univ. 47 (2005), 1–28; _01_Kishi.pdf; 1035.pdf (223 K), escholarship@OUDIR; MR2198857 (2007f:11125) (summary) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1161.11396 (from the introduction) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[29] 内藤浩忠 : Spiegelungssatz について, 第 3 回北陸数論研究集会報告集 [Proceedings of the 3rd Workshop on Number Theory] (金沢大学サテライトプラザ, 2004) (平林幹人, 野村明人, 木村巌, 菅野孝史編), 2005, pp. 25–32.

[30] 岸康弘 : Spiegelungssatz の構成的アプローチ, 第 3 回北陸数論研究集会報告集 [Proceedings of the 3rd Workshop on Number Theory] (金沢大学サテライトプラザ, 2004) (平林幹人, 野村明人, 木村巌, 菅野孝史編), 2005, pp. 33–49.

[31] 山村健 : Spiegelungssatz に関する文献第 3 回北陸数論研究集会報告集 [Proceedings of the 3rd Workshop on Number Theory] (金沢大学サテライトプラザ, 2004) (平林幹人, 野村明人, 木村巌, 菅野孝史編), 2005, pp. 120–123.

[32] J. Oh : A note on the first layers of ℤ_{p}-extensions, Commun. Korean Math. Soc. 24 (2009), no. 1, 1–4; [PDF]; DOI : 10.4134/CKMS.2009.24.1.001 [DBSHCJ_2009_v24n1_1.pdf] MR2488805 (2010a:11212) (reviewed by Q. Ji) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1168.11324 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[33] É. Fouvry and J. Klüners : On the Spiegelungssatz for the 4-rank, Algebra & Number Theory 4 (2010), no. 5, 493– 508; DOI: 10.2140/ant.2010.4.493 [ant-v4-n5-p01-s.pdf]; MR2679097 (2011i:11159) (reviewed by M. Horie) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1225.11140 (reviewed by B. Sodaïgui) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[34] L. Habsieger and E. Royer : Spiegelungssatz: a combinatorial proof for the 4-rank, Int. J. Number Theory 7 (2011), no. 8, 2157–2170; DOI: 10.1142/S1793042111005106 [PDF (222 KB)].